OpenStudy (anonymous):

How many solutions exist for the system y = 2x + 5 and y = -0.2x - 4?

OpenStudy (anonymous):

Many solutions !!!

OpenStudy (anonymous):

Look : y = 2x + 5 y = -0.2x - 4 So we can say : 2x+5=-2/10x-4 and then have : 2x+2/10x=-9 2/2x=-9 Now find x . !